### v3

@@ Abstract
COVID-19 pandemic has brought to the fore epidemiological models which, though describing a wealth of behaviors, have previously received little attention in the signal processing literature. In this work, a generalized time-varying autoregressive model is considered, encompassing, but not reducing to, a state-of-the-art model of viral epidemics propagation. The time-varying parameter of this model is estimated via the minimization of a penalized likelihood estimator. A major challenge is that the estimation accuracy strongly depends on hyperparameters fine-tuning. Without available ground truth, hyperparameters are selected by minimizing specifically designed data-driven oracles, used as proxy for the estimation error. Focusing on the time-varying autoregressive Poisson model, the Stein’s Unbiased Risk Estimate formalism is generalized to construct asymptotically unbiased risk estimators based on the derivation of an original autoregressive counterpart of Stein’s lemma. The accuracy of these oracles and of the resulting estimates are assessed through intensive Monte Carlo simulations on synthetic data. Then, elaborating on recent epidemiological models, a novel weekly scaled Poisson model is proposed, better accounting for intrinsic variability of the contamination while being robust to reporting errors. Finally, the overall data-driven procedure is particularized to the estimation of COVID-19 reproduction number demonstrating its ability to yield very consistent estimates.
@@ 1 Introduction
Context. Inverse problems are ubiquitous in signal and image processing [40, 32, 82, 52], with a wealth of domains of application as diverse as nonlinear physics [16, 60], astronomy [20], hyperspectral imaging [7], tomography [69], cardiology [4] and epidemiology [59]. A general inverse problem consists in estimating underlying quantities of interest from direct or indirect observations. The intricate measurement processes necessary to obtain physical1 observations can be the source of several difficulties in estimating the quantities of interest, among which: acquisition performed in a transformed domain requiring backward transformation to access the quantity of interest [69], linear or nonlinear deformation of observations through the measurement process [8, 64], and corruption by stochastic perturbations, either stemming from the physical [^1] Physical is to be understood in a very broad sense, encompassing biological, epidemiological as well as human data. phenomenon at stake or from the measurement device [32]. Mathematically, without loss of generality, the quantity of interest and observations can be represented by real-valued vectors $\boldsymbol{\mathsf{X}}\in\mathbb{R}^{T}$ and $\boldsymbol{\mathsf{Y}}\in\mathbb{R}^{S}$ respectively, with $T,S\in\mathbb{N}^{*}$, and a generic inverse problem writes
@@
$\boldsymbol{\lambda}^{\dagger}\in\underset{\boldsymbol{\lambda}\in\Lambda}{\mathrm{Argmin}}\,\,\left\lVert\widehat{\boldsymbol{\mathsf{X}}}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})-\overline{\boldsymbol{\mathsf{X}}}\right\rVert^{2}_{2}.$ (3)
Though, in practice, the ground truth is not available, and hence the estimation error cannot be evaluated. To tackle this issue, several classes of hyperparameter tuning strategies have been developed during the past decades. For additive Gaussian noise models, leading to quadratic data-fidelity terms, regularized with quadratic penalizations, the L-curve criterion consists in selecting the regularization level by close inspection of the plot of the residual against the regularity of the solution as the regularization parameter is varied [37]. Not only the L-curve method is restricted to Gaussian models under Tikhonov regularization, but also it has been shown to be inaccurate when the targeted ground truth is very smooth [36] and to behave inconsistently as the size of the problem increases [78]. An alternative way is to model the hyperparameters as random variables and to estimate them via hierarchical Bayesian techniques [68, 32]. Such strategies come at the price of an extra complexity, as they require to specify the a priori hyperparameters distribution, and a significant computational cost, as Monte Carlo sampling is often necessary. Recent deep learning methods take advantage of large training databases to learn the hyperparameters of variational estimators [5, 55, 30]. While being very accurate as soon as enough annotated data are available, these methods are not adapted to tackle fundamentally new problems for which the generation of training database would either be too costly or even not possible due to lack of available expert knowledge. A widely used class of unsupervised methods consists in the construction of an oracle $\mathcal{O}$ not depending explicitly on the ground truth, which can thus be evaluated in practice, and whose minimization yields an approximation of the optimal hyperparameter in terms of the true error (3) defined as
@@
Model (5) is very reminiscent of Problem (1), the unknown quantity $\boldsymbol{\mathsf{X}}$ being $(\mathsf{R}_{1},\ldots,\mathsf{R}_{T})$, the role of A being played by the linear operator:
$(\mathsf{R}_{1},\ldots,\mathsf{R}_{T})\mapsto\left(\Phi_{1}(\boldsymbol{\mathsf{Z}})\mathsf{R}_{1},\ldots,\Phi_{T}(\boldsymbol{\mathsf{Z}})\mathsf{R}_{T}\right)$ (6)
and the degradation $\mathcal{B}$ consisting in data-dependent Poisson noise. Elaborating on this formal resemblance, the variational framework (2) has been fruitfully leveraged to design COVID-19 instantaneous reproduction number $\mathsf{R}_{t}$ estimators [1, 59]. Up to now, the fine-tuning of the regularization parameters of these variational estimators has been done manually, based on expert knowledge, which not only impairs the analysis of huge amount of data, but also might reflect subjective biases of the users. To derive a fully data-driven hyperparameter selection strategy, the main challenge lies in the design of an adapted oracle $\mathcal{O}$, for example of a Stein estimator. Indeed, meticulous comparison of Models (1) and (5) shows that the autoregressive nature of the epidemiological model (5), which induce a dependency of the linear operator $(\mathsf{R}_{1},\ldots,\mathsf{R}_{T})\mapsto\left(\Phi_{1}(\boldsymbol{\mathsf{Z}})\mathsf{R}_{1},\ldots,\Phi_{T}(\boldsymbol{\mathsf{Z}})\mathsf{R}_{T}\right)$ in the observation vector $\boldsymbol{\mathsf{Z}}$, definitively excludes direct use of generalized Stein Unbiased Risk Estimators, requiring A to be statistically independent of $\boldsymbol{\mathsf{Z}}$, if not deterministic [71, 23, 60, 61]. Developing a novel Stein paradigm, adapted to autoregressive models, is a challenging, though crucial, step toward the design of fully data-driven, hence objective, strategies for reproduction number estimation.
@@ 2.1 Observation model
The present work considers a generalized time-varying autoregressive model encompassing both the standard autoregressive model of finite order and the epidemiological model introduced in [17] while generalizing both of them in several directions. The first major originality is that the model is externally controlled by an unknown deterministic time-varying reproduction coefficient containing the information of interest. This setup drastically contrasts both with the standard time series framework, mostly interest in predicting the behavior of the observations rather than inferring the coefficients [18], and with random coefficient autoregressive models in which the time-varying coefficients are random variable [38, 63]. Second, observations are no longer restricted to follow an independent Gaussian distribution with constant variance, but instead any distribution with prescribed mean, possibly depending on additional time-varying parameters. Importantly the noise is not necessary additive but might be data-dependent or multiplicative. And finally, the memory term $\Psi_{t}$ is not necessarily a linear function, it is only assumed to be causal, that is depending only on past observations, and smooth.
@@
where $\mathcal{G}\left(\alpha,\mathsf{U}\right)$ refers to the Gamma distribution of shape parameter $\alpha$ and scale parameter $\mathsf{U}$ [^3] By convention, whatever $\alpha>0$, if $\mathsf{U}\leq 0$, $\mathcal{P}(\mathsf{U})$ and $\mathcal{G}\left(\alpha,\mathsf{U}\right)$ are Dirac distributions, i.e., $\mathsf{Y}_{t}=0$ deterministically. Both the Gaussian and Gamma distributions are well-suited to described continuous quantities. The seminal additive Gaussian noise model is much used in biomedical signal analysis, for example to analyze electroencephalogram (EEG) recordings [84, 12], and in communications [83, 77], while the multiplicative Gamma noise model has been used in financial econometrics [45]. Data-dependent Poisson noise is mostly used for count time series in which the observations can take only discrete, typically integer, values [63], with applications in epidemiology [81, 17] and in climate studies, for example to analyze changepoints in the North Atlantic tropical cyclone record [67]. Figure 1 shows on the top row generalized time-varying autoregressive data under Model (7) with respectively additive Gaussian noise 1(a), data-dependent Poisson noise 1(b), and multiplicative Gamma noise 1(c) with prescribed time-varying reproduction coefficient, displayed in dark blue on the bottom row. For all three examples, the function $\psi$ shaping the memory term corresponds to a discretized Gamma distribution of mean $6.6$ and standard deviation $3.5$ truncated at $\tau=25$, mimicking the serial interval function of COVID-19 [9, 66].
@@ 2.2 Variational estimators
where $\boldsymbol{\lambda}=(\lambda_{1},\ldots,\lambda_{L})\in\Lambda$ is a vector of regularization parameters, balancing the overall regularization level as well as the relative importance of the different constraints encoded in the penalization. Commonly used regularization terms are composite [44, 64, 65] and are expressed as
@@
where for each $\ell\in\{1,\ldots,L\}$, $\textbf{L}_{\ell}$ is a linear operator, $q_{\ell}$ a positive exponent, and $\lambda_{\ell}\geq 0$ is a regularization parameter balancing the importance of the $\ell$th constraint with respect to the other constraints in (24) and to the data-fidelity term of Equation (23). Each term of the functional enforces a specific constraint, hence enabling to take into account several regularity and sparsity properties simultaneously. For example, when choosing the discrete Laplacian $\textbf{L}=\textbf{D}_{2}$, $q_{\ell}=1$, favors sparsity of the second order temporal derivative, and hence results in piecewise linear estimates, while $q=2$ yields smooth estimates. The penalized likelihood strategy sketched in Equation (23) is highly flexible and adapts to a large collection of noise models and constraints, hence, by favoring a priori behavior, it has the ability to provide consistent and accurate regularized estimates. The excellent performance of variational estimators comes at the price of a cautious fine-tuning of the regularization parameters associated to each term of the penalization. Not only is this task very tough to perform manually but also, and more importantly, in practice ground truth is not available, and it is necessary to resort to data-driven oracles to approach optimal hyperparameters.
@@
Figure 1: Generalized time-varying autoregressive data and Maximum Likelihood reproduction coefficient estimates. The processes $\mathsf{Y}_{t}$, displayed as solid black curves on top row, follow Model (7) with piecewise linear reproduction coefficient $\mathsf{X}_{t}$, displayed as dark blue curve on the bottom row; linear memory functions (8) with horizon $\tau=25$; initialized at $\mathsf{Y}_{0}=10^{4}$ and run for $T=300$ steps. The associated memory terms $\Psi_{t}(\boldsymbol{\mathsf{Y}})$ are displayed as dashed black curves on the top row. Blue areas on the bottom row corresponds to $\mathsf{X}_{t}\geq 1$ inducing exponential growth of $\mathsf{Y}_{t}$. The Maximum likelihood estimates (13) of the underlying reproduction coefficient under each model are displayed as light blue curves on the bottom row. (a) Additive Gaussian noise (10) with constant standard deviation $\alpha_{t}\equiv 10^{5}$. (b) Poisson degradation (11), with constant scale parameter $\alpha_{t}\equiv 10^{3}$. (c) Gamma degradation (12), with constant shape parameter $\alpha_{t}\equiv 20$.
@@ 3.1 General framework
where $\mathbb{E}_{\boldsymbol{\mathsf{Y}}}$ denotes the expectation over realizations of $\boldsymbol{\mathsf{Y}}$. For inverse problems of the form (1) with an ill-conditioned or non-injective operator A, the estimation risk is potentially numerically instable; an alternative is to shift the emphasis on the reconstruction error and to consider the prediction risk
@@
Both the estimation and prediction risks depends on the ground truth $\overline{\boldsymbol{\mathsf{X}}}$, which in practice is not available. The purpose of this section is thus to devise oracles for the quality an estimate $\widehat{\boldsymbol{\mathsf{X}}}$, which are independent of the unknown ground truth, and whose minimization provides approximately optimal hyperparameters, where optimal is to be understood as reaching low estimation or prediction risk.
@@ 3.2 A novel autoregressive Poisson lemma
The remaining of the paper focuses on the Poisson model (11), which is commonly used for modeling the pathogen spread during epidemics, with the aim of applying the developed tools to the estimation of COVID-19 reproduction number from real infection counts in Section 5.
@@
Let $\Theta:\mathbb{R}^{T}\rightarrow\mathbb{R}$, the real-valued function defined on $\mathbb{N}^{T}$ as
$\boldsymbol{\mathsf{k}}\mapsto\Theta(\alpha_{1}\mathsf{k}_{1},\ldots,\alpha_{T}\mathsf{k}_{T})\overline{\mathsf{X}}_{t}\Psi_{t}(\alpha_{1}\mathsf{k}_{1},\ldots,\alpha_{t-1}\mathsf{k}_{t-1})$
is summable with respect to the autoregressive Poisson distribution obtained plugging the generalized time-varying autoregression model (7) into the Poisson distribution (11)
@@
It is worth insisting on the fact that, due to the dependency of the memory functions in the past observations, the components of $\boldsymbol{\mathsf{Y}}$ are not independent Poisson random variables. This is visible in Equation (27) where, because $\Psi_{s}(\boldsymbol{\mathsf{Y}})$ depends on $\mathsf{Y}_{1},\ldots,\mathsf{Y}_{s-1}$, the right-hand side is not a separable product of independent Poisson distribution and cannot be reframed as such. This is precisely this major difference with standard inverse problems of the form (1) which impairs the application of the standard Poisson counterpart of Stein’s lemma [51, 47, 48]. [^5] This remark on data-dependent Poisson noise would be true as well for the additive Gaussian and multiplicative Gamma noises of Example 2.

### v4

@@ Abstract
COVID-19 pandemic has brought to the fore epidemiological models which, though describing a wealth of behaviors, have previously received little attention in signal processing literature. In this work, a generalized time-varying autoregressive model is considered, encompassing, but not reducing to, a state-of-the-art model of viral epidemics propagation. The time-varying parameter of this model is estimated via the minimization of a penalized likelihood estimator. A major challenge is that the estimation accuracy strongly depends on hyperparameters fine-tuning. Without available ground truth, hyperparameters are selected by minimizing specifically designed data-driven oracles, used as proxy for the estimation error. Focusing on the time-varying autoregressive Poisson model, Stein’s Unbiased Risk Estimate formalism is generalized to construct asymptotically unbiased risk estimators based on the derivation of an original autoregressive counterpart of Stein’s lemma. The accuracy of these oracles and of the resulting estimates are assessed through intensive Monte Carlo simulations on synthetic data. Then, elaborating on recent epidemiological models, a novel weekly scaled Poisson model is proposed, better accounting for intrinsic variability of the contaminations while being robust to reporting errors. Finally, the data-driven procedure is particularized to the estimation of COVID-19 reproduction number from weekly infection counts demonstrating its ability to tackle real-world applications.
@@ 1 Introduction
Context. Inverse problems are ubiquitous in signal and image processing [40, 32, 83, 52], with a wealth of domains of application as diverse as nonlinear physics [16, 60], astronomy [20], hyperspectral imaging [7], tomography [69], cardiology [4] and epidemiology [59]. A general inverse problem consists in estimating underlying quantities of interest from direct or indirect observations. The intricate measurement processes necessary to obtain physical1 observations can be the source of several difficulties in estimating the quantities of interest, among which: acquisition performed in a transformed domain requiring backward transformation to access the quantity of interest [69], linear or nonlinear deformation of observations through the measurement process [8, 64], and corruption by stochastic perturbations, either stemming from the physical [^1] Physical is to be understood in a very broad sense, encompassing biological, epidemiological as well as human data. phenomenon at stake or from the measurement device [32]. Mathematically, without loss of generality, the quantities of interest and observations can be represented by real-valued vectors $\boldsymbol{\mathsf{X}}\in\mathbb{R}^{T}$ and $\boldsymbol{\mathsf{Y}}\in\mathbb{R}^{S}$ respectively, with $T,S\in\mathbb{N}^{*}$, and a generic inverse problem writes
@@
$\boldsymbol{\lambda}^{\dagger}\in\underset{\boldsymbol{\lambda}\in\Lambda}{\mathrm{Argmin}}\,\,\left\lVert\widehat{\boldsymbol{\mathsf{X}}}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})-\overline{\boldsymbol{\mathsf{X}}}\right\rVert^{2}_{2}$ (3)
where $\overline{\boldsymbol{\mathsf{X}}}$ denotes the ground truth quantity of interest. Though, in practice, the lack of ground truth impairs evaluation of the estimation error. To tackle this issue, several classes of hyperparameter tuning strategies have been developed during the past decades. For additive Gaussian noise models, leading to quadratic data-fidelity terms, regularized with quadratic penalizations, the L-curve criterion consists in selecting the regularization level by close inspection of the plot of the residual against the regularity of the solution as the regularization parameter is varied [37]. Not only the L-curve method is restricted to Gaussian models under Tikhonov regularization, but also it has been shown to be inaccurate when the targeted ground truth is very smooth [36] and to behave inconsistently as the size of the problem increases [79]. An alternative way is to model the hyperparameters as random variables and to estimate them via hierarchical Bayesian techniques [68, 32]. Such strategies come at the price of an extra complexity, as they require to specify the a priori hyperparameters distribution, and a significant computational cost, as Monte Carlo sampling is often necessary. Recent deep learning methods take advantage of large training databases to learn the hyperparameters of variational estimators [5, 55, 30]. While being very accurate as soon as enough annotated data are available, these methods are not adapted to tackle fundamentally new problems for which the generation of training database would either be too costly or even not possible due to lack of available expert knowledge. A widely used class of unsupervised methods consists in the construction of an oracle $\mathcal{O}$ not depending explicitly on the ground truth, which can thus be evaluated in practice, and whose minimization yields an approximation of the optimal hyperparameter $\boldsymbol{\lambda}^{\dagger}$ defined as
@@
Model (5) is very reminiscent of Problem (1), the unknown quantity $\boldsymbol{\mathsf{X}}$ being $(\mathsf{R}_{1},\ldots,\mathsf{R}_{T})$, the role of A being played by the linear operator $(\mathsf{R}_{1},\ldots,\mathsf{R}_{T})\mapsto\left(\Phi_{1}(\boldsymbol{\mathsf{Z}})\mathsf{R}_{1},\ldots,\Phi_{T}(\boldsymbol{\mathsf{Z}})\mathsf{R}_{T}\right)$ and the degradation $\mathcal{B}$ consisting in data-dependent Poisson noise. Elaborating on this formal resemblance, the variational framework (2) has been fruitfully leveraged to design COVID-19 effective reproduction number $\mathsf{R}_{t}$ estimators [1, 59]. Up to now, the fine-tuning of the regularization parameters of these variational estimators has been done manually, based on expert knowledge, which not only impairs the analysis of huge amount of data, but also might reflect subjective biases of the users. To derive a fully data-driven hyperparameter selection strategy, the main challenge lies in the design of an adapted oracle $\mathcal{O}$, for example of a Stein estimator. Indeed, meticulous comparison of Models (1) and (5) shows that the autoregressive nature of the epidemiological model (5), which induce a dependency of the linear operator $(\mathsf{R}_{1},\ldots,\mathsf{R}_{T})\mapsto\left(\Phi_{1}(\boldsymbol{\mathsf{Z}})\mathsf{R}_{1},\ldots,\Phi_{T}(\boldsymbol{\mathsf{Z}})\mathsf{R}_{T}\right)$ in the observation vector $\boldsymbol{\mathsf{Z}}$, definitively excludes direct use of generalized Stein Unbiased Risk Estimators, requiring A to be statistically independent of $\boldsymbol{\mathsf{Z}}$, if not deterministic [71, 23, 60, 61]. Developing a novel Stein paradigm, adapted to autoregressive models, is a challenging, though crucial, step toward the design of fully data-driven, hence objective, strategies for reproduction number estimation.
@@ 2.1 Observation model
The present work considers a generalized time-varying autoregressive model encompassing both the standard autoregressive model of finite order and the epidemiological model introduced in [17] while generalizing both of them in several directions. The first major originality is that the model is externally controlled by an unknown deterministic time-varying reproduction coefficient containing the information of interest. This setup drastically contrasts both with the standard time series framework, mostly interested in predicting the behavior of the observations rather than inferring the coefficients [18], and with random coefficient autoregressive models in which the time-varying coefficients are random variables [38, 63]. Second, observations are no longer restricted to follow an independent Gaussian distribution with constant variance, but instead any distribution with prescribed mean, possibly depending on additional time-varying parameters. Importantly the noise is not necessary additive but might be data-dependent or multiplicative. And finally, the memory term $\Psi_{t}$ is not necessarily a linear function, it is only assumed to be causal, that is depending only on past observations, and smooth.
@@
where $\mathcal{G}\left(\alpha,\mathsf{U}\right)$ refers to the Gamma distribution of shape parameter $\alpha$ and scale parameter $\mathsf{U}$. [^3] By convention, whatever $\alpha>0$, if $\mathsf{U}\leq 0$, $\mathcal{P}(\mathsf{U})$ and $\mathcal{G}\left(\alpha,\mathsf{U}\right)$ are Dirac distributions, i.e., $\mathsf{Y}_{t}=0$ deterministically. Both the Gaussian and Gamma distributions are well-suited to described continuous quantities. The seminal additive Gaussian noise model is much used in biomedical signal analysis, for example to analyze electroencephalogram (EEG) recordings [85, 12], and in communications [84, 78], while the multiplicative Gamma noise model has been used in financial econometrics [45]. Data-dependent Poisson noise is mostly used for count time series in which the observations can take only discrete, typically integer, values [63], with applications in epidemiology [82, 17] and in climate studies, for example to analyze changepoints in the North Atlantic tropical cyclone record [67]. Figure 1 shows on the top row generalized time-varying autoregressive data under Model (6) with respectively additive Gaussian noise 1(a), data-dependent Poisson noise 1(b), and multiplicative Gamma noise 1(c) with prescribed time-varying reproduction coefficient, displayed in dark blue on the bottom row. For all three examples, the function $\psi$ shaping the memory term corresponds to a discretized Gamma distribution of mean $6.6$ and standard deviation $3.5$ truncated at $\tau=25$, mimicking the serial interval function of COVID-19 [9, 66].
@@ 2.2 Variational estimators
where $\boldsymbol{\lambda}=(\lambda_{1},\ldots,\lambda_{L})\in\Lambda\subseteq\mathbb{R}_{+}^{L}$ is a vector of regularization parameters, balancing the overall regularization level as well as the relative importance of the different constraints encoded in the penalization. Commonly used regularization terms are composite [44, 64, 65] and are expressed as
@@
[…]
@@
Figure 1: Generalized time-varying autoregressive data and Maximum Likelihood reproduction coefficient estimates. The processes $\mathsf{Y}_{t}$, displayed as solid black curves on top row, follow Model (6) with piecewise linear reproduction coefficient $\mathsf{X}_{t}$, displayed as dark blue curve on the bottom row; linear memory functions (7) with horizon $\tau=25$; initialized at $\mathsf{Y}_{0}=10^{4}$ and run for $T=300$ steps. The associated memory terms $\Psi_{t}(\boldsymbol{\mathsf{Y}})$ are displayed as dashed black curves on the top row. Blue areas on the bottom row correspond to $\mathsf{X}_{t}\geq 1$ inducing exponential growth of $\mathsf{Y}_{t}$. The Maximum likelihood estimates (12) of the underlying reproduction coefficient under each model are displayed as light blue curves on the bottom row. (a) Additive Gaussian noise (9) with constant standard deviation $\alpha_{t}\equiv 10^{5}$. (b) Poisson degradation (10), with constant scale parameter $\alpha_{t}\equiv 10^{3}$. (c) Gamma degradation (11), with constant shape parameter $\alpha_{t}\equiv 20$.
@@ 3.1 General framework
where $\mathbb{E}_{\boldsymbol{\mathsf{Y}}}$ denotes the expectation over realizations of $\boldsymbol{\mathsf{Y}}$ and $\overline{\boldsymbol{\mathsf{X}}}$ the ground truth. For inverse problems of the form (1) with an ill-conditioned or non-injective operator A, the estimation risk is potentially numerically instable; an alternative is to shift the emphasis on the reconstruction error and to consider the prediction risk
@@
Both the estimation and prediction risks depends on the ground truth $\overline{\boldsymbol{\mathsf{X}}}$, which in practice is not available. The purpose of this section is thus to devise oracles for the quality of an estimate $\widehat{\boldsymbol{\mathsf{X}}}$, which are independent of the unknown ground truth, and whose minimization provides approximately optimal hyperparameters, where optimal is to be understood as reaching low estimation (24) or prediction (25) risk.
@@ 3.2 A novel autoregressive Poisson lemma
The remaining of the paper focuses on the Poisson model (10), which is commonly used for modeling the pathogen spread during epidemics, with the aim of applying the developed tools to the estimation of COVID-19 reproduction number from real infection counts in Section 5. It is worth noting that similar proof mechanisms could be applied to extend the analysis to Gaussian (9) and Gamma (11) noise distributions.
@@
Let $\Theta:\mathbb{R}^{T}\rightarrow\mathbb{R}$, the real-valued function defined on $\mathbb{N}^{T}$ as $\boldsymbol{\mathsf{k}}\mapsto\Theta(\alpha_{1}\mathsf{k}_{1},\ldots,\alpha_{T}\mathsf{k}_{T})\overline{\mathsf{X}}_{t}\Psi_{t}(\alpha_{1}\mathsf{k}_{1},\ldots,\alpha_{t-1}\mathsf{k}_{t-1})$ is summable with respect to the autoregressive Poisson distribution obtained plugging the generalized time-varying autoregressive model (6) into the Poisson distribution (10)
@@
It is worth insisting on the fact that, due to the dependency of the memory functions in the past observations, the components of $\boldsymbol{\mathsf{Y}}$ are not independent Poisson random variables. This is visible in Equation (26) where, because $\Psi_{s}(\boldsymbol{\mathsf{Y}})$ depends on $\mathsf{Y}_{1},\ldots,\mathsf{Y}_{s-1}$, the right-hand side is not a separable product of independent Poisson distribution and cannot be reframed as such. This is precisely this major difference with standard inverse problems of the form (1) which impairs the application of the standard Poisson counterpart of Stein’s lemma [51, 47, 48]. [^5] This remark on data-dependent Poisson noise applies as well for the additive Gaussian and multiplicative Gamma noises of Example 2.